# Droplet coalescence kinetics: Coalescence mechanisms and thermodynamic non-equilibrium effects under isothermal and non-isothermal conditions

### v1

@@ Abstract
Droplet coalescence is a multiscale phenomenon driven by multiple thermodynamic non-equilibrium (TNE) effects. This study investigates the droplet coalescence mechanisms and the interplay between various TNE effects under isothermal and non-isothermal conditions kinetically. The main findings include: (1) Coalescence initiation and cut-through mechanisms: In non-isothermal conditions, the temperature rise caused by the release of latent heat during phase transition slightly increases the surface tension gradient (driving force) near the contact point of the two droplets, while significantly enhancing the pressure gradient (resistance). This results in a significantly prolonged coalescence initiation time compared to the isothermal case. The additional temperature field effect increases the spatial complexity of both the flow field and droplet morphology. Regarding the cut-through mechanism, in both cases, pressure extends the liquid-vapor interface in opposite directions, promoting the growth of the liquid bridge. (2) TNE effects: Latent heat-induced temperature rise significantly refrains the TNE intensity in thermal case. Before and after droplet contact, non-equilibrium quantities driven by the temperature gradient ($\bm{\Delta}^{\ast}_{3,1}$ and $\bm{\Delta}^{\ast}_{3}$) and those driven by the velocity gradient ($\bm{\Delta}^{\ast}_{2}$ and $\bm{\Delta}^{\ast}_{4,2}$), alternate in dominating the coalescence process. This competition and interplay result in a more complex spatial and spatiotemporal evolution of TNE effects compared to the isothermal case, where the TNE intensity is primarily dominated by $\bm{\Delta}^{\ast}_{2}$ and $\bm{\Delta}^{\ast}_{4,2}$. (3) Entropy production mechanisms: In the non-isothermal case, entropy production is contributed not only by $\bm{\Delta}^{\ast}_{2}$ but also by $\bm{\Delta}^{\ast}_{3,1}$, with the former being the dominant contributor. The temperature field reduces the entropy production rate, while extends its duration, and increases the total entropy production. Before and after droplet cut-through, the entropy production rates from $\bm{\Delta}^{\ast}_{2}$ and $\bm{\Delta}^{\ast}_{3,1}$ alternate in dominating the total entropy production during coalescence. Further investigation into the initial droplet distance $r_{d}$ effect on coalescence revealed that, $r_{d}$ increases droplet cut-through time $t_{c}$, the time when total non-equilibrium reaches its maximum $t_{D\rm max}$, and the time when total entropy production rate reaches its maximum $t_{\dot{S}\rm prmax}$ in a power-law manner: $(t_{c},t_{D\rm max},t_{\dot{S}\rm prmax})\sim{r_{d}}^{\alpha}$. This research provides kinetic insights into droplet coalescence, offering mesoscopic support for dynamic, cross-scale regulation and multifunctional integration of coalescence processes in industrial applications.
@@ I Introduction
Droplet coalescence is a widespread phenomenon in fields such as nature, biomedical sciences, microfluidics, engine combustion, multi-medium vehicles, and photolithographyKaminski and Garstecki (2017); Moragues et al. (2023); Li et al. (2021a); Sitte and Mastorakos (2019); Jiaxing et al. (2023); Yang et al. (2023); Yang, Wei, and Guo (2023); Chen et al. (2021). In the field of microfluidics, droplet-based technologies involving deformation, fragmentation, coalescence, and collision, hold great potential for applications in medical diagnostics, food science, drug development, and artificial biomimetic systems manufacturingKaminski and Garstecki (2017); Moragues et al. (2023); Chen et al. (2021). Droplets, serve as multifunctional carriers for controlled reactions, mixing, and analysis, play a key role in these processesKaminski and Garstecki (2017); Moragues et al. (2023); Shan et al. (2023). In combustion systems with dense spray devices, intense interactions between droplets-including atomization, collision, rupture, coalescence, and oscillation-greatly affect fuel heat and mass transfer, combustion efficiency, and combustion stability Li et al. (2021a); Sitte and Mastorakos (2019). In explosion dynamics, studying droplet deformation, collision, and coalescence under shock waves at various Mach numbers provides crucial insights for mitigating shockwave energy and reducing shock damage Jiaxing et al. (2023).
@@
The main difference between thermal and isothermal multiphase flow systems is the heat generation and conduction during phase separation, resulting in lower interfacial stresses and localized temperature changes due to latent heat release, which leads to new local mechanical equilibrium Gan et al. (2011). Under non-isothermal conditions, the variety of driving mechanisms increases significantly, and the competition and coupling between these mechanisms intensify, resulting in a more complex droplet coalescence process. Specifically, due to the sensitivity of surface tension to temperature, different surface tensions arise on the droplet surface. The resulting surface tension gradient drives the droplet to flow along the temperature gradient, a phenomenon known as the Marangoni effect. This effect significantly influences coalescence dynamics Sun et al. (2021); Vrij and Overbeek (1968); Perumanath et al. (2019); Geri et al. (2017). The local temperature difference resulting from the gradient also alters the liquid’s viscosity, enhancing both heat and mass transfer, and the coupling between them. These changes modify the internal convection patterns and interface behavior of the droplet, subsequently affecting its deformation velocity and coalescence dynamics He, He, and Zhang (2024); Sun et al. (2021). Currently, the kinetic study of droplet collision, coalescence, and fragmentation under non-isothermal conditions remains limited. Non-isothermal conditions, however, are prevalent in combustion systems Li et al. (2021a); Sitte and Mastorakos (2019); He, He, and Zhang (2024); Sun et al. (2021), explosion dynamics Jiaxing et al. (2023), nuclear reactors Liang et al. (2022), and microfluidics Dixit et al. (2010); Davanlou and Kumar (2015). For instance, in microfluidics, the thermal Marangoni effect is often used to control the droplet coalescence process Dixit et al. (2010); Davanlou and Kumar (2015). This highlights the importance of comparative studies on the driving mechanisms and underlying kinetic behaviors of droplet coalescence under isothermal and non-isothermal conditions.
@@
i) Neglect of temperature field and compressibility effects: Most studies currently rely on isothermal, incompressible models, neglecting the impact of temperature fields and compressible convection on flow behavior. As a result, these models fail to fully capture the intricate interactions between droplets in real-world multiphase flows.
ii) Oversight of non-equilibrium behaviors and effects: While most studies focus on the evolution of conserved quantities—such as mass, momentum, and energy—alongside slow-varying variables, they often overlook the evolution of non-conserved, fast-varying variables like thermodynamic non-equilibrium (TNE) and hydrodynamic non-equilibrium (HNE) behaviors. These non-equilibrium phenomena are closely tied to small-scale structures and fast modes, such as liquid bridges, droplet boundary layers, vortices, droplet fragments, splash phenomena, and surface fluctuations. In regions near these structures, discrete effects become pronounced, challenging the reliability of the methods discussed above. Furthermore, during rapid collisions and coalescence, the onset of fast modes impedes the system’s return to thermodynamic equilibrium in a timely manner, leading to strong non-equilibrium effects that further undermine the validity of these approaches. On the mesoscopic scale, HNE and TNE arise from the deviation of the distribution function from equilibrium, highlighting system complexity. On the macroscopic scale, these non-equilibrium effects are intricately coupled with macroscopic quantities and exert a fundamental influence on the overall behavior of the system.
iii) Limitations in data analysis and information extraction: In the droplet coalescence process, various driving mechanisms—including TNE effects, surface tension, temperature gradients, and velocity gradients—interact to produce complex, multi-scale spatiotemporal structures. To achieve a deeper understanding of these processes, a multi-perspective analysis is essential.
iv) Lack of systematic comparison of entropy production mechanisms: Entropy production mechanisms are pivotal in describing the thermodynamic behavior during droplet coalescence, particularly with regard to energy conversion and dissipation in non-equilibrium states. Despite their importance, existing studies lack a systematic comparison of entropy production mechanisms, especially in terms of the contributions of different driving forces—such as surface tension, heat flux, and viscous stress—to the entropy production rate. This gap limits our understanding of entropy production and non-equilibrium effects during droplet coalescence and hinders the broader applicability of existing models in complex systems.
@@
Recent studies have emphasized the importance of investigating TNE effects and behaviors. In scenarios where TNE occurs, neglecting its effects can lead to inaccurate estimations of viscosity, heat flux, and ultimately result in imprecise calculations of macroscopic quantities Gan et al. (2022); Zhang et al. (2019a); Lin and Luo (2019). In the field of complex multiphase flows, such as phase separation Gan et al. (2015); Zhang et al. (2019b), bubble/droplet coalescence Zhang et al. (2020); Sun et al. (2022, 2024), and laser-bubble/droplet interactions Zhang et al. (2023), TNE effects offer significant advantages in characterizing the evolutionary features of scaled systems. For example, during liquid-vapor phase separation, the maximum points of the total TNE intensity and entropy production rate can be used to distinguish between the spinodal decomposition and domain growth stages Gan et al. (2015); Zhang et al. (2019b). The non-organized momentum flux component in the coalescing direction aids in identifying stages in bubble coalescence Sun et al. (2022). Additionally, the intensity of NOMF can provide a standard for distinguishing different stages of the droplet collision process and for differentiating between types of droplet collisions Zhang et al. (2020). In fluid instability systems, by detecting TNE features from different perspectives, a more detailed physical structure of the interface can be provided, which can then be used for the physical identification and tracking technology design of different interfaces within the systemGan et al. (2019); Lai et al. (2016); Chen et al. (2024a, b); Song et al. (2024).
@@ II The model
where $f_{ki}^{eq}$ is the discrete form of the local equilibrium distribution function Watari and Tsutahara (2003), with $k$ representing the $k$-th group of particle velocities $\mathbf{v}_{k}$, and $i$ the direction of $\mathbf{v}_{k}$. The external force term, $I_{ki}=-[A+\mathbf{B}\cdot(\mathbf{v}_{ki}-\mathbf{u})+(C+C_{q})(\mathbf{v}_{ki}-\mathbf{u})^{2}]f_{ki}^{eq}$, is introduced using a mean-field approximation Gonnella, Lamura, and Sofonea (2007); Gan et al. (2022) to account for the contribution of intermolecular interaction forces to the rate of change of the distribution function. Here, $A=-2(C+C_{q})T$ ensures mass conservation. $\mathbf{B}=\frac{1}{\rho T}\bm{\nabla}\cdot\left[(P^{\mathrm{CS}}-\rho T)\mathbf{I}+\bm{\Lambda}\right]$with $\bm{\Lambda}=K\bm{\nabla}\rho\bm{\nabla}\rho-K(\rho\bm{\nabla}^{2}\rho+\frac{1}{2}|\bm{\nabla}\rho|^{2})\mathbf{I}-[\rho T\bm{\nabla}\rho\cdot\bm{\nabla}(K/T)]\mathbf{I}$ representing the density gradient contribution to the pressure tensor, $\mathbf{I}$ is the unit tensor and $K$ is the surface tension coefficient. The term $C=\frac{1}{{2\rho{T^{2}}}}\{\left({{P^{\mathrm{CS}}}-\rho T}\right)\bm{\nabla}\cdot\mathbf{u}+{\bm{\Lambda}}:\bm{\nabla}{\bf{u}}+a{\rho^{2}}\bm{\nabla}\cdot\mathbf{u}-K\left[{\frac{1}{2}\bm{\nabla}\rho\cdot\bm{\nabla}\rho\bm{\nabla}\cdot\mathbf{u}+\rho\bm{\nabla}\rho\cdot\bm{\nabla}\left({\bm{\nabla}\cdot\mathbf{u}}\right)+\bm{\nabla}\rho\cdot\bm{\nabla}{\bf{u}}\cdot\bm{\nabla}\rho}\right]\}$ represents a partial contribution to energy. The term $C_{q}=\frac{1}{\rho T^{2}}\bm{\nabla}\cdot(q\rho T\bm{\nabla}T)$ accounts for additional heat conduction within the system. The Carnahan-Starling equation of state (EOS) is given by $P^{\mathrm{CS}}=\rho T\frac{1+\eta+\eta^{2}-\eta^{3}}{(1-\eta)^{3}}-a\rho^{2}$, with $\eta=\frac{b\rho}{4}$, and $a$ and $b$ being the attraction and repulsion parameters, respectively. Here, $\rho$, $\mathbf{u}$, and $T$ represent the local density, velocity, and temperature, respectively.
@@
$\bm{\Delta}_{m,n}=\mathbf{M}_{m,n}(f_{ki}-f_{ki}^{eq})=\sum\limits_{ki}\left(\frac{1}{2}\right)^{1-\delta_{m,n}}(f_{ki}-f_{ki}^{eq})(\mathbf{v}_{ki}\cdot\mathbf{v}_{ki})^{(m-n)/2}\mathbf{v}_{ki}^{n},$ (2)
@@
Among the TNE measures, $\bm{\Delta}_{2}^{*}$ represents the non-organized momentum flux (NOMF) or stress flux, with its first-order analytical form: $\bm{\Delta}_{2}^{*(1)}=-\rho T\tau\left[\nabla\mathbf{u}+(\nabla\mathbf{u})^{T}-\mathbf{I}\nabla\cdot\mathbf{u}\right]$. The term $\bm{\Delta}_{3,1}^{*}$ represents the non-organized energy flux (NOEF) or heat flux, with its first-order analytical expression: $\bm{\Delta}_{3,1}^{*(1)}=-2\rho T\tau\nabla T$. The term $\bm{\Delta}_{3}^{*}$ represents the viscous stress flux, and $\bm{\Delta}_{4,2}^{*}$ represents the flux of heat flux. To provide a comprehensive description of the system’s non-equilibrium intensity, DBM adopts the non-equilibrium intensity vector: $\mathbf{S}_{\text{TNE}}=\{|\bm{\Delta}_{2}^{*}|,|\bm{\Delta}_{3,1}^{*}|,|\bm{\Delta}_{3}^{*}|,|\bm{\Delta}_{4,2}^{*}|,\dots,|\bm{\Delta}_{m,n}^{*}|,{D^{*}}\}$.
For a dissipative system, the entropy production rate is closely related to TNE quantities:
@@
The total entropy production rate reads $\dot{S}_{\rm pr}=\dot{S}_{\rm NOMF}+\dot{S}_{\rm NOEF}$.
III Numerical simulations and analysis
All parameters in the simulations are dimensionless. The reference variables are chosen as the reference density $\rho_{\text{r}}$, reference temperature $T_{\text{r}}$, and the global size of the domain $L_{\text{r}}$, to facilitate the non-dimensionalization of the actual physical quantities. The choice of reference quantities is arbitrary; for instance, the critical density $\rho_{c}$ and critical temperature $T_{c}$ can be selected as reference variables. The relationship between the dimensionless parameters in DBM and the physical ones are:
$\hat{\rho}=\frac{\rho}{{{\rho_{\text{r}}}}},\quad\hat{T}=\frac{T}{{{T_{\text{r}}}}},\quad\hat{r}_{\alpha}=\frac{r_{\alpha}}{{{L_{\text{r}}}}},$ (5)
$\left({\hat{t},\hat{\tau}}\right)=\frac{{\left({t,\tau}\right)}}{{{L_{\text{r}}}/\sqrt{R{T_{\text{r}}}}}},\quad\hat{u}_{\alpha}=\frac{u_{\alpha}}{\sqrt{R{T_{\text{r}}}}},\quad{\hat{c}_{p}}=\frac{{{c_{p}}}}{R},$ (6)
$\hat{\mu}=\frac{\mu}{{{\rho_{\text{r}}}{L_{\text{r}}}\sqrt{R{T_{\text{r}}}}}},\quad\hat{\kappa}=\frac{\kappa}{{{\rho_{\text{r}}}{L_{\text{r}}}R\sqrt{R{T_{\text{r}}}}}},\quad\hat{\sigma}=\frac{\sigma}{{{\rho_{\text{r}}}R{T_{\text{r}}}{L_{\text{r}}}}},$ (7)

### v3

@@ Abstract
Droplet coalescence is a multiscale phenomenon driven by multiple thermodynamic non-equilibrium (TNE) effects. This study investigates the coalescence mechanisms and the complex interplay between various TNE effects in both isothermal and non-isothermal cases kinetically. The main findings include: (1) Coalescence initiation and cut-through mechanisms: In the non-isothermal case, latent heat release results in a temperature rise that slightly increases the surface tension gradient (driving force) near the contact point, while substantially enhancing the pressure gradient (resistance). This leads to a significantly prolonged initiation of coalescence compared to the isothermal case. The additional temperature field effects introduce greater spatial complexity to both the flow field and morphology. For the cut-through mechanism, in both cases, pressure extends the liquid-vapor interface in opposite directions, promoting liquid bridge growth. (2) TNE effects: The latent heat-induced temperature rise in the thermal case considerably dampens the intensity of TNE effects. Before and after droplet contact, TNE quantities driven by the temperature gradient ($\bm{\Delta}^{\ast}_{3,1}$ and $\bm{\Delta}^{\ast}_{3}$) and those driven by the velocity gradient ($\bm{\Delta}^{\ast}_{2}$ and $\bm{\Delta}^{\ast}_{4,2}$), alternate in dominating the coalescence process. This alternating dominance result in a more complex spatiotemporal evolution of TNE effects compared to the isothermal case, where the TNE intensity is dominated by $\bm{\Delta}^{\ast}_{2}$ and $\bm{\Delta}^{\ast}_{4,2}$. (3) Entropy production mechanisms: In the non-isothermal case, entropy production is contributed by both $\bm{\Delta}^{\ast}_{2}$ and $\bm{\Delta}^{\ast}_{3,1}$, with the former being the dominant contributor. The temperature field reduces the entropy production rate, but extends its duration, thereby increasing the total entropy production. The entropy production rates from $\bm{\Delta}^{\ast}_{2}$ and $\bm{\Delta}^{\ast}_{3,1}$ alternate before and after the droplet cut-through. Further investigation into effects of the initial droplet distance $r_{d}$ revealed that it increases droplet cut-through time $t_{c}$, the time when total TNE reaches its maximum $t_{D\rm max}$, and the time when total entropy production rate reaches its maximum $t_{\dot{S}\rm prmax}$ in a power-law manner: $(t_{c},t_{D\rm max},t_{\dot{S}\rm prmax})\sim{r_{d}}^{\alpha}$. This research provides kinetic insights into droplet coalescence, offering mesoscopic support for dynamic, cross-scale regulation and multifunctional integration of coalescence processes in industrial applications.
@@ I Introduction
Droplet coalescence is a widespread phenomenon in fields such as nature, biomedical sciences, microfluidics, engine combustion, multi-medium vehicles, and photolithographyKaminski and Garstecki (2017); Moragues et al. (2023); Li et al. (2021a); Sitte and Mastorakos (2019); Jiaxing et al. (2023); Yang, Wei, and Guo (2023); Yang et al. (2023); Chen et al. (2021). In the field of microfluidics, droplet-based technologies involving deformation, fragmentation, coalescence, and collision, hold great potential for applications in medical diagnostics, food science, drug development, and artificial biomimetic systems manufacturingKaminski and Garstecki (2017); Moragues et al. (2023); Chen et al. (2021). Droplets, serve as multifunctional carriers for controlled reactions, mixing, and analysis, play a key role in these processesKaminski and Garstecki (2017); Moragues et al. (2023); Shan et al. (2023). In combustion systems with dense spray devices, intense interactions between droplets-including atomization, collision, rupture, coalescence, and oscillation-greatly affect fuel heat and mass transfer, combustion efficiency, and combustion stabilityLi et al. (2021a); Sitte and Mastorakos (2019). In explosion dynamics, studying droplet deformation, collision, and coalescence under shock waves at various Mach numbers provides crucial insights for mitigating shockwave energy and reducing shock damageJiaxing et al. (2023).
@@
The main difference between thermal and isothermal multiphase flow systems is the heat generation and conduction during phase separation, resulting in lower interfacial stresses and localized temperature changes due to latent heat release, which leads to new local mechanical equilibriumGan et al. (2011). Under non-isothermal conditions, the variety of driving mechanisms increases significantly, and the competition and coupling between these mechanisms intensify, resulting in a more complex droplet coalescence process. Specifically, because surface tension varies sensitively with temperature, a gradient forms across the droplet surface, propelling the droplet along the temperature gradient—a process known as the Marangoni effect. This effect significantly influences coalescence dynamicsSun et al. (2021); Vrij and Overbeek (1968); Perumanath et al. (2019); Geri et al. (2017). The local temperature difference resulting from the gradient also alters the liquid’s viscosity, enhancing both heat and mass transfer, and the coupling between them. These changes modify the internal convection patterns and interface behavior of the droplet, subsequently affecting its deformation velocity and coalescence dynamicsHe, He, and Zhang (2024); Sun et al. (2021). Currently, the kinetic study of droplet collision, coalescence, and fragmentation under non-isothermal conditions remains limited. Non-isothermal conditions, however, are prevalent in combustion systemsLi et al. (2021a); Sitte and Mastorakos (2019); He, He, and Zhang (2024); Sun et al. (2021), explosion dynamicsJiaxing et al. (2023), nuclear reactorsLiang et al. (2022), and microfluidicsDixit et al. (2010); Davanlou and Kumar (2015). For instance, in microfluidics, the thermal Marangoni effect is often used to control the droplet coalescence processDixit et al. (2010); Davanlou and Kumar (2015). This highlights the importance of comparative studies on the driving mechanisms and underlying kinetic behaviors of droplet coalescence under isothermal and non-isothermal conditions.
@@
(i) Neglect of temperature field and compressibility effects: Most studies currently rely on isothermal, incompressible models, neglecting the impact of temperature fields and compressible convection on flow behavior. As a result, these models fail to fully capture the intricate interactions between droplets in real-world multiphase flows.
(ii) Oversight of non-equilibrium behaviors and effects: While most studies focus on the evolution of conserved quantities—such as mass, momentum, and energy—alongside slow-varying variables, they often overlook the evolution of non-conserved, fast-varying variables like thermodynamic non-equilibrium (TNE) and hydrodynamic non-equilibrium (HNE) behaviors. These non-equilibrium phenomena are closely tied to small-scale structures and fast modes, such as liquid bridges, droplet boundary layers, vortices, droplet fragments, splash phenomena, and surface fluctuations. In regions near these structures, discrete effects become pronounced, challenging the reliability of the methods discussed above. Furthermore, during rapid collisions and coalescence, the onset of fast modes impedes the system’s return to thermodynamic equilibrium in a timely manner, leading to strong non-equilibrium effects that further undermine the validity of these approaches. On the mesoscopic scale, HNE and TNE arise from the deviation of the distribution function from equilibrium, highlighting system complexity. On the macroscopic scale, these non-equilibrium effects are intricately coupled with macroscopic quantities and exert a fundamental influence on the overall behavior of the system.
(iii) Limitations in data analysis and information extraction: In the droplet coalescence process, various driving mechanisms—including TNE effects, surface tension, temperature gradients, and velocity gradients—interact to produce complex, multi-scale spatiotemporal structures. To achieve a deeper understanding of these processes, a multi-perspective analysis is essential.
(iv) Lack of systematic comparison of entropy production mechanisms: Entropy production mechanisms are pivotal in describing the thermodynamic behavior during droplet coalescence, particularly with regard to energy conversion and dissipation in non-equilibrium states. Despite their importance, existing studies lack a systematic comparison of entropy production mechanisms, especially in terms of the contributions of different driving forces—such as surface tension, heat flux, and viscous stress—to the entropy production rate. This gap limits our understanding of entropy production and non-equilibrium effects during droplet coalescence and hinders the broader applicability of existing models in complex systems.
@@
Recent studies have emphasized the importance of investigating TNE effects and behaviors. In scenarios where TNE occurs, neglecting its effects can lead to inaccurate estimations of viscosity, heat flux, and ultimately result in imprecise calculations of macroscopic quantitiesGan et al. (2022); Zhang et al. (2019a); Lin and Luo (2019). In the field of complex multiphase flows, such as phase separationGan et al. (2015); Zhang et al. (2019b), bubble/droplet coalescenceZhang et al. (2020); Sun et al. (2022, 2024), and laser-bubble/droplet interactionsZhang et al. (2023), TNE effects offer significant advantages in characterizing the evolutionary features of scaled systems. For example, during liquid-vapor phase separation, the maximum points of the total TNE intensity and entropy production rate can be used to distinguish between the spinodal decomposition and domain growth stagesGan et al. (2015); Zhang et al. (2019b). The non-organized momentum flux component in the coalescing direction aids in identifying stages in bubble coalescenceSun et al. (2022). Additionally, the intensity of non-organized momentum flux (NOMF) can provide a standard for distinguishing different stages of the droplet collision process and for differentiating between types of droplet collisionsZhang et al. (2020). In fluid instability systems, by detecting TNE features from different perspectives, a more detailed physical structure of the interface can be provided, which can then be used for the physical identification and tracking technology design of different interfaces within the systemGan et al. (2019); Lai et al. (2016); Chen et al. (2024a, b); Song et al. (2024).
@@ II The model
where $f_{ki}^{eq}$ is the discrete form of the local equilibrium distribution function Watari and Tsutahara (2003), with $k$ representing the $k$-th group of particle velocities $\mathbf{v}_{ki}$, and $i$ the direction of $\mathbf{v}_{ki}$. The external force term, $I_{ki}=-[A+\mathbf{B}\cdot(\mathbf{v}_{ki}-\mathbf{u})+(C+C_{q})(\mathbf{v}_{ki}-\mathbf{u})^{2}]f_{ki}^{eq}$, is introduced using a mean-field approximation Gan et al. (2022); Gonnella, Lamura, and Sofonea (2007) to account for the contribution of intermolecular interaction forces to the rate of change of the distribution function. Here, $A=-2(C+C_{q})T$ ensures mass conservation. $\mathbf{B}=\frac{1}{\rho T}\bm{\nabla}\cdot\left[(P-\rho T)\mathbf{I}+\bm{\Lambda}\right]$with $\bm{\Lambda}=K\bm{\nabla}\rho\bm{\nabla}\rho-K(\rho\bm{\nabla}^{2}\rho+\frac{1}{2}|\bm{\nabla}\rho|^{2})\mathbf{I}-[\rho T\bm{\nabla}\rho\cdot\bm{\nabla}(K/T)]\mathbf{I}$ representing the density gradient contribution to the pressure tensor, $\mathbf{I}$ is the unit tensor and $K$ is the surface tension coefficient. The term $C=\frac{1}{{2\rho{T^{2}}}}\{\left({{P}-\rho T}\right)\bm{\nabla}\cdot\mathbf{u}+{\bm{\Lambda}}:\bm{\nabla}{\bf{u}}+a{\rho^{2}}\bm{\nabla}\cdot\mathbf{u}-K\left[{\frac{1}{2}\bm{\nabla}\rho\cdot\bm{\nabla}\rho\bm{\nabla}\cdot\mathbf{u}+\rho\bm{\nabla}\rho\cdot\bm{\nabla}\left({\bm{\nabla}\cdot\mathbf{u}}\right)+\bm{\nabla}\rho\cdot\bm{\nabla}{\bf{u}}\cdot\bm{\nabla}\rho}\right]\}$ represents a partial contribution to energy. The term $C_{q}=\frac{1}{\rho T^{2}}\bm{\nabla}\cdot(q\rho T\bm{\nabla}T)$ accounts for additional heat conduction within the system. The Carnahan-Starling equation of state (EOS) is given by $P=\rho T\frac{1+\eta+\eta^{2}-\eta^{3}}{(1-\eta)^{3}}-a\rho^{2}$, with $\eta=\frac{b\rho}{4}$, and $a$ and $b$ being the attraction and repulsion parameters, respectively. Here, $\rho$, $\mathbf{u}$, and $T$ represent the local density, velocity, and temperature, respectively.
@@
$\bm{\Delta}_{m,n}$ $=\mathbf{M}_{m,n}(f_{ki}-f_{ki}^{eq})$ (2)
$=\sum\limits_{ki}$ $\left(\frac{1}{2}\right)^{1-\delta_{m,n}}(f_{ki}-f_{ki}^{eq})(\mathbf{v}_{ki}\cdot\mathbf{v}_{ki})^{\frac{m-n}{2}}\mathbf{v}_{ki}^{n},$
@@
Among the TNE measures, $\bm{\Delta}_{2}^{*}$ represents the NOMF or viscous stress, with its first-order analytical form: $\bm{\Delta}_{2}^{*(1)}=-\rho T\tau\left[\bm{\nabla}\mathbf{u}+(\bm{\nabla}\mathbf{u})^{T}-\mathbf{I}\bm{\nabla}\cdot\mathbf{u}\right]$. The term $\bm{\Delta}_{3,1}^{*}$ represents the non-organized energy flux (NOEF) or heat flux, with its first-order analytical expression: $\bm{\Delta}_{3,1}^{*(1)}=-2\rho T\tau\bm{\nabla}T$. The term $\bm{\Delta}_{3}^{*}$ represents the viscous stress flux, and $\bm{\Delta}_{4,2}^{*}$ represents the flux of heat flux. To provide a comprehensive description of the system’s non-equilibrium intensity, DBM adopts the non-equilibrium intensity vector: $\mathbf{S}_{\text{TNE}}=\{|\bm{\Delta}_{2}^{*}|,|\bm{\Delta}_{3,1}^{*}|,|\bm{\Delta}_{3}^{*}|,|\bm{\Delta}_{4,2}^{*}|,\dots,|\bm{\Delta}_{m,n}^{*}|,{D^{*}}\}$, with ${D{{}^{\ast}=}}\sum_{m,n}|\bm{\Delta}_{m,n}^{\ast}|$ the total TNE intensity. The definitions, components, and their physical meanings are displayed in the Appendix.
Taking the moments of Eq. (1) with the collision invariant vector ($1$, ${\mathbf{v}_{ki}}$, ${{{v}_{ki}^{2}}/2}$), the generalized hydrodynamic equations (GHEs) for non-ideal fluids considering surface tension effect are obtained:
${\frac{{\partial\rho}}{{\partial t}}}+\bm{\nabla}\cdot(\rho\mathbf{u})=0,$ (3)
${\frac{{\partial(\rho\mathbf{u})}}{{\partial t}}}+\bm{\nabla}\cdot(\rho\mathbf{uu}+P\mathbf{I})+\bm{\nabla}\cdot(\bm{\Lambda}+\bm{\Delta}_{2}^{\ast})=0,$ (4)
with ${e_{T}}=\rho T-a{\rho^{2}}+{{K{{\left|{\bm{\nabla}\rho}\right|}^{2}}}/2+\rho u^{2}/2}$ the total energy density. Recovering the GHEs represents only one aspect of the physical capabilities of the DBM. In fact, the full physical description provided by DBM corresponds to the extended hydrodynamic equations Gan et al. (2022), which encompass not only the evolution equations for conserved moments, but also for non-conserved moments closely associated with thermodynamic and hydrodynamic non-equilibrium effects. The significance of these extended equations increases notably as the discrete effects or non-equilibrium intensity becomes stronger.
Both types of non-equilibrium effects, NOMF and NOEF contribute to the entropy production rate,
$\dot{S}_{\rm pr}=\dot{S}_{\rm NOMF}+\dot{S}_{\rm NOEF},$ (6)
where
@@
III Simulations and results
Droplet coalescence is a multiscale phenomenon driven by the interplay of multiple physical mechanisms, particularly the competing TNE effects. This study systematically explores these mechanisms by examining: (i) slowly varying macroscopic variables, (ii) interface dynamics, (iii) rapidly varying non-equilibrium quantities, and (iv) entropy production rates under both isothermal and non-isothermal conditions.
All parameters in the manuscript and simulations are dimensionless. The reference variables can be chosen as the reference density $\rho_{\text{r}}$, reference temperature $T_{\text{r}}$, and the global size of the domain $L_{\text{r}}$. The choice of the reference quantity is arbitrary; for instance, the critical density $\rho_{c}$ and the critical temperature $T_{c}$ can be selected as reference variables, calculated from the non-ideal equation of state. The relationship between the parameters in DBM and the physical ones are as follows:
$\hat{\rho}=\frac{\rho}{{{\rho_{\text{r}}}}},\quad\hat{T}=\frac{T}{{{T_{\text{r}}}}},\quad\hat{r}_{\alpha}=\frac{r_{\alpha}}{{{L_{\text{r}}}}},\quad\hat{u}_{\alpha}=\frac{u_{\alpha}}{\sqrt{R{T_{\text{r}}}}},$ (9)
$\left({\hat{t},\hat{\tau}}\right)=\frac{{\left({t,\tau}\right)}}{{{L_{\text{r}}}/\sqrt{R{T_{\text{r}}}}}},\quad{\hat{c}_{p}}=\frac{{{c_{p}}}}{R},\quad\hat{\sigma}=\frac{\sigma}{{{\rho_{\text{r}}}R{T_{\text{r}}}{L_{\text{r}}}}},$ (10)
$\hat{\mu}=\frac{\mu}{{{\rho_{\text{r}}}{L_{\text{r}}}\sqrt{R{T_{\text{r}}}}}},\quad\hat{\kappa}=\frac{\kappa}{{{\rho_{\text{r}}}{L_{\text{r}}}R\sqrt{R{T_{\text{r}}}}}},$ (11)